OpenStudy (anonymous):

The difference of two integers is 23. If the larger integer is divided by the smaller one, the quotient is 2 with a remainder of 1. What are the two integers? A. 40 and 63 B. 45 and 22 C. 38 and 61 D. 30 and 53

OpenStudy (ajprincess):

Let the largest integer be x and the smallest be y. x-y=23 If the larger integer is divided by the smaller one, the quotient is 2 with a remainder of 1 means the larger integer is 1 more than twice of the smaller integer. that is x=2y+1 Nw solve the simultaneous eqns to find the numbers.

OpenStudy (anonymous):

i dont know to do simultaneous eqns

OpenStudy (ajprincess):

x-y=23-(1) Substitute x in (1) with 2y+1. So nw (1) becomes 2y+1-y=23-(2) Nw can u simplify (2)?

OpenStudy (anonymous):

im so confused

OpenStudy (ajprincess):

can u simplify 2y+1-y=23?

OpenStudy (anonymous):

no

OpenStudy (ajprincess):